OpenStudy (anonymous):

Find the oblique asymptote. (x^2-4x-5)/(x-3)

myininaya (myininaya):

3| 1 -4 -5 | 3 -3 | _____________ 1 -1 | -8 so we have \[\frac{x^2-4x-5}{x-3}=x-1+\frac{-8}{x-3}\] so y=x-1 is the slant asymptote